# DEER: A Comprehensive and Reliable Benchmark for Deep-Research Expert Reports

### v1

@@ C.2 Construction Procedure and Evidence Mapping
To verify the validity of the drafted dimensions and detailed criteria, this study conducted cross-validation with 10 independent experts representing 6 domains (Computer Science, Artificial Intelligence, Biology, Chemistry, Mathematics, Psychology). Two experts per domain reviewed the suitability of each item, focusing on factors considered necessary in light of domain practices, such as whether each dimension and criterion is actually a meaningful and necessary evaluation standard in that domain, whether the definition and scope of application are excessively ambiguous, and whether it unnecessarily overlaps with other criteria. Based on this, they made binary (pass/fail) evaluations for each criterion. The final evaluation framework included only the criteria that passed this cross-validation process.
@@ Appendix D Foundations of Deep Research Report Evaluation Taxonomy
Request Completeness aligns with guideline-level completeness checks that evaluate whether all required components of a report are present according to prescribed inclusion mandates across authoritative standards Bastian2021PRISMA; Boutron2010CONSORT; ISO25010; WWC_Handbook; Percie2020ARRIVE; GRI_Standards; STROBE_2007. Scope Boundary reflects boundary-setting requirements that assess whether a report explicitly specifies its operative limits—such as eligibility conditions, assumptions, and constraints—within the structural fields defined in major evaluative frameworks Bastian2021PRISMA; NIST_AI_RMF; AHA2024Standards; AEA2024Data; OSF_Prereg; NASA_SE_Handbook; SEC_Reg_SK; SPIRIT_2013. Value Validity corresponds to coherence and utility standards that evaluate whether the report’s core conclusions not only align logically with the evidence base but also deliver sufficient specificity and practical feasibility to comprehensively address the user’s inquiry Guyatt2013GRADE; APA_JARS2025; NBER2024; IES_SEER; OECD_Policy_Eval.
Evidence Validity
@@
Numeric Accuracy reflects evaluative requirements that check whether quantitative statements in a report correspond to verifiable computations, prespecified analytic procedures, and reproducible numerical evidence as established in major methodological frameworks Guyatt2013GRADE; JCGM_GUM; ACM_Badging; IUPAC_Green; IUPAP_Red; Econometric_Soc; AERA_APA_NCME_Standards; CODATA_2022; OECD_Quality; IFRS_Standards; ISO30414; IPCC_2019. Logical Support aligns with reasoning-assessment criteria that evaluate whether inferential steps are explicitly grounded in evidence, free of unsupported leaps, and consistent with theoretical or mathematical rigor EMS_Code2025; APA_GoodPractices; AERA2006Standards; AMS_Handbook; Fed_Rules_Evidence; COREQ_2007.
Structure Consistency
@@
Introduction corresponds to structural-orientation requirements that assess whether a report’s opening section presents its scope and analytic pathway in accordance with prescribed organizational fields in established reporting and specification frameworks Bastian2021PRISMA; Boutron2010CONSORT; ISO29148_2018; IFRS_Integrated; CARE_2013. Body aligns with structural-progression criteria that evaluate whether the main sections develop the promised analytical sequence without omission or drift relative to the ordered components defined in recognized guideline structures Bastian2021PRISMA; Boutron2010CONSORT; ISO29148_2018; CF_Conventions; CARE_2013. Conclusion reflects closure-consistency requirements that examine whether final statements integrate preceding evidence without introducing unsupported expansions, consistent with the conclusion-governance rules embedded in major evaluative standards Bastian2021PRISMA; Boutron2010CONSORT; CARE_2013. Section-Level corresponds to intra- and inter-section coherence checks that assess alignment, ordering, and non-duplication of content in accordance with structured specification and reporting templates in authoritative frameworks Bastian2021PRISMA; ISO29148_2018; IFRS_Integrated.
Narration & Style
@@
Report Form corresponds to format-governance requirements that evaluate whether a report adheres to prescribed structural conventions for scientific communication as codified in established editorial and reporting standards Bastian2021PRISMA; Boutron2010CONSORT; ICMJE2025; Leipzig_Rules; SIAM_Style; IPA_Handbook; APSA_Style; ASA_Style. Writing Quality aligns with language-precision criteria that assess clarity, specificity, and terminological consistency according to recognized guidelines for accurate and unbiased scholarly expression Boutron2010CONSORT; ICMJE2025; Chicago_Style; Blacks_Law; HBR_Guide. Paragraph reflects cohesion-assessment rules that examine whether individual paragraphs follow a coherent internal logic—anchoring topic statements, supporting evidence, and transitional structure—in line with authoritative reporting and specification frameworks Bastian2021PRISMA; ISO29148_2018; The_Bluebook; ASA_Style. Reader Friendliness corresponds to accessibility-oriented requirements that evaluate whether the narrative facilitates comprehension through appropriate signaling, explanatory devices, and presentation practices as outlined in major editorial and reporting guidelines Bastian2021PRISMA; ICMJE2025; HBR_Guide; GRI_Standards.
@@
Sensitive Issues corresponds to ethical-screening requirements that evaluate whether sensitive domains are addressed with neutrality, respect, and responsible language in accordance with established editorial and reporting ethics standards ICMJE2025; Bastian2021PRISMA; OHA_Principles; OAH_Standards; BPS_Code; Belmont_Report; APSA_Ethics. Safety & Impact aligns with harm-assessment provisions that examine whether potential adverse effects, misuse risks, or disproportionate impacts are identified and mitigated under recognized guidelines for responsible scientific communication ICMJE2025; Guyatt2013GRADE; NRC_Prudent; Helsinki_Decl; ESA_Code. Perspective Balance reflects fairness-evaluation criteria that assess whether multiple viewpoints and counterpositions are represented without undue bias following principles of impartiality articulated in authoritative publication and reporting frameworks ICMJE2025; Bastian2021PRISMA; BPA_GoodPractice; AAP_Code; CFA_Code; ABA_Model_Rules; AAPOR_Code.
@@
Source Support aligns with evidence-provision requirements that evaluate whether claims are accompanied by verifiable supporting sources in accordance with established evidentiary and reporting guidelines Bastian2021PRISMA; FORCE11_Data; NBER2024; ISSCR2025; GIPS_2020. Information corresponds to content-adequacy criteria that assess whether a report supplies all necessary factual and contextual material needed to justify its reasoning under recognized completeness and transparency standards Bastian2021PRISMA; ISO25010; EQUATOR_Network; LSA_Ethics2024; JOC_Guidelines; USGS_FSP; GRI_Standards. Citations align with source-attribution rules that examine whether in-text citations are provided at appropriate argumentative locations following authoritative norms for evidentiary traceability ICMJE2025; Wilkinson2016FAIR; IEEE2025Ethics; ACS2021Ethical; The_Bluebook. References are evaluated to determine whether the number of actually used references is appropriate and whether they collectively form a well-documented, traceable, and accessible evidence base StanfordAI2025; FORCE11_Data; DA_RT_Principles.
@@
Claim Factuality reflects evidence-verification provisions that assess whether factual assertions in a report correspond to verifiable sources and documented evidence traces within established evaluative frameworks Wilkinson2016FAIR; APS2023Conduct; NIST_AI_RMF; PDG_Review; NIST_DLMF; SEC_Reg_SK. Citation Validity corresponds to source-justification checks that evaluate whether cited references substantively support the claims they accompany according to recognized standards for evidential accountability Wilkinson2016FAIR; ASA_Ethics; AHA2024Standards; Austin_Principles. Reference Accuracy aligns with source-quality and provenance requirements that examine whether referenced materials reliably and transparently support the arguments for which they are cited, in line with authoritative data-governance and reporting criteria Wilkinson2016FAIR; NBER2024; FORCE11_Data; APA_JARS2025; IFRS_Standards. Reference Quality evaluates whether the set of sources consists of reproducible, trustworthy, and methodologically sound information sources, ensuring that the report’s evidentiary foundation is grounded in reliable materials ACS2021Ethical; IEEE2025Ethics; SEP_Policy; Helsinki_Decl; Belmont_Report. Reference Diversity evaluates whether the evidence base avoids excessive concentration on a small number of sources and instead maintains a balanced level of source diversity, reducing the risk of narrow, biased, or selectively framed arguments EQUATOR_Network; LSA_Ethics2024; AAPOR_Code.
@@
Domain (80 listings) Reference
Artificial Intelligence (8) NIST_AI_RMF; StanfordAI2025; ACL2024Guidelines; NeurIPS2025Guidelines; ICML2025Instructions; ICLR2024Guidelines; JMLR2024AuthorGuide; Springer2025MLJournal
Biology (4) Bastian2021PRISMA; Boutron2010CONSORT; ISSCR2025; Percie2020ARRIVE
Business (4) GRI_Standards; HBR_Guide; ISO30414; IFRS_Integrated
Chemistry (4) ACS2021Ethical; IUPAC_Green; JOC_Guidelines; NRC_Prudent
Computer Science (4) ISO25010; ACM_Badging; ISO29148_2018; Wilkinson2016FAIR
Earth & Env. Science (4) IPCC_2019; USGS_FSP; CF_Conventions; ESA_Code
Economics (4) NBER2024; AEA2024Data; Econometric_Soc; OECD_Quality
Education (4) WWC_Handbook; AERA2006Standards; IES_SEER; AERA_APA_NCME_Standards
Engineering (4) IEEE_Std2025; IEEE2025Ethics; NASA_SE_Handbook; ISO29148_2018
Finance (4) IFRS_Standards; CFA_Code; GIPS_2020; SEC_Reg_SK
History (4) AHA2024Standards; Chicago_Style; OHA_Principles; OAH_Standards
Law (4) The_Bluebook; Blacks_Law; Fed_Rules_Evidence; ABA_Model_Rules
Linguistics (4) LSA_Ethics2024; Leipzig_Rules; Austin_Principles; IPA_Handbook
Mathematics (4) EMS_Code2025; AMS_Handbook; SIAM_Style; NIST_DLMF
Medicine (4) Helsinki_Decl; STROBE_2007; CARE_2013; SPIRIT_2013
Philosophy (4) APA_GoodPractices; SEP_Policy; BPA_GoodPractice; AAP_Code
Physics (4) APS2023Conduct; IUPAP_Red; PDG_Review; CODATA_2022
Political Science (4) DA_RT_Principles; APSA_Style; OECD_Policy_Eval; APSA_Ethics
Psychology (4) APA_JARS2025; OSF_Prereg; BPS_Code; AERA_APA_NCME_Standards
Sociology (4) AAPOR_Code; COREQ_2007; Belmont_Report; ASA_Style
@@ Appendix E Rubric Structure
Dimension / Criterion Element / Rubric description
2. Evidence Validity └ 2.2 Logical Support 2.2.5 Element All major claims follow logically from the previously presented facts, data, interpretations, and assumptions, without skipped steps or unsupported leaps. [2.2.5.1 Coverage] All claims that require logical support are explicitly linked to the relevant facts, data, and interpretations (including the key evidence specified in the Expert Evaluation Guidance, EG), and no central claim is left without the necessary supporting evidence. [2.2.5.2 Quality] Each major claim is explained through clear, well-structured reasoning that shows how the underlying facts, data, and interpretations lead to that conclusion, making the logical connection explicit and easy for an expert reader to follow.
Table 11: Example Coverage/Quality rubric factors for a reasoning-focused element under 2. Evidence Validity / 2.2 Logical Support.
@@ E.1 Hierarchical Levels
Level 2: Detailed Criteria (26)
Detailed criteria that decompose each dimension into finer evaluation units; 2–5 detailed criteria are defined under each dimension. This forms an intermediate-level unit that specifies which aspects to examine for each dimension.
@@
Thus, with 7 dimensions, 26 detailed criteria, and 40 elements as the framework, each element is divided into Coverage/Quality factors to constitute a total of 130 specific check items. Table 11 shows an example of one element and its corresponding Coverage/Quality factors in the 2. Evidence Validity / 2.2 Logical Support dimension.

### v2

@@ C.2 Construction Procedure and Evidence Mapping
To verify the validity of the drafted dimensions and sub-dimensions, this study conducted cross-validation with 10 independent experts representing 6 domains (Computer Science, Artificial Intelligence, Biology, Chemistry, Mathematics, Psychology). Two experts per domain reviewed the suitability of each item, focusing on factors considered necessary in light of domain practices, such as whether each dimension and criterion is actually a meaningful and necessary evaluation standard in that domain, whether the definition and scope of application are excessively ambiguous, and whether it unnecessarily overlaps with other criteria. Based on this, they made binary (pass/fail) evaluations for each criterion. The final evaluation framework included only the criteria that passed this cross-validation process.
@@ Appendix D Foundations of Deep Research Report Evaluation Taxonomy
Completeness aligns with guideline-level completeness checks that evaluate whether all required components of a report are present according to prescribed inclusion mandates across authoritative standards Bastian and Moher (2021); Boutron et al. (2010); International Organization for Standardization (2011); What Works Clearinghouse (2022); Percie du Sert et al. (2020); Global Reporting Initiative (2023); von Elm and others (2007). Scope reflects boundary-setting requirements that assess whether a report explicitly specifies its operative limits—such as eligibility conditions, assumptions, and constraints—within the structural fields defined in major evaluative frameworks Bastian and Moher (2021); National Institute of Standards and Technology (2023); American Historical Association (2024); American Economic Association (2024); Center for Open Science (2024); NASA (2016); U.S. Securities and Exchange Commission (2024); Chan and others (2013). Helpfulness corresponds to coherence and utility standards that evaluate whether the report’s core conclusions not only align logically with the evidence base but also deliver sufficient specificity and practical feasibility to comprehensively address the user’s inquiry Guyatt et al. (2013); American Psychological Association (2025); Chang et al. (2024); Institute of Education Sciences (2022); OECD (2020).
Analytical Soundness
@@
Quantification reflects evaluative requirements that check whether quantitative statements in a report correspond to verifiable computations, prespecified analytic procedures, and reproducible numerical evidence as established in major methodological frameworks Guyatt et al. (2013); Joint Committee for Guides in Metrology (2008); Association for Computing Machinery (2025); International Union of Pure and Applied Chemistry (2007); International Union of Pure and Applied Physics (2010); The Econometric Society (2024); American Educational Research Association et al. (2014); Mohr and others (2024); OECD (2011); International Accounting Standards Board (2024); International Organization for Standardization (2018); Intergovernmental Panel on Climate Change (2019). Reasoning aligns with reasoning-assessment criteria that evaluate whether inferential steps are explicitly grounded in evidence, free of unsupported leaps, and consistent with theoretical or mathematical rigor European Mathematical Society (2025); American Philosophical Association (2024); American Educational Research Association (2006); American Mathematical Society (2022); U.S. Congress (2024); Tong et al. (2007).
Structural Coherence
@@
Introduction corresponds to structural-orientation requirements that assess whether a report’s opening section presents its scope and analytic pathway in accordance with prescribed organizational fields in established reporting and specification frameworks Bastian and Moher (2021); Boutron et al. (2010); IEEE Computer Society (2018); IFRS Foundation (2021); Gagnier and others (2013). Body aligns with structural-progression criteria that evaluate whether the main sections develop the promised analytical sequence without omission or drift relative to the ordered components defined in recognized guideline structures Bastian and Moher (2021); Boutron et al. (2010); IEEE Computer Society (2018); Eaton and others (2024); Gagnier and others (2013). Conclusion reflects closure-consistency requirements that examine whether final statements integrate preceding evidence without introducing unsupported expansions, consistent with the conclusion-governance rules embedded in major evaluative standards Bastian and Moher (2021); Boutron et al. (2010); Gagnier and others (2013). Section-Level corresponds to intra- and inter-section coherence checks that assess alignment, ordering, and non-duplication of content in accordance with structured specification and reporting templates in authoritative frameworks Bastian and Moher (2021); IEEE Computer Society (2018); IFRS Foundation (2021).
Format & Style
@@
Report Format corresponds to format-governance requirements that evaluate whether a report adheres to prescribed structural conventions for scientific communication as codified in established editorial and reporting standards Bastian and Moher (2021); Boutron et al. (2010); International Committee of Medical Journal Editors (2025); Comrie, B. and Haspelmath, M. and Bickel, B. (2015); Society for Industrial and Applied Mathematics (2024); International Phonetic Association (1999); American Political Science Association (2018); American Sociological Association (2022). Writing Quality aligns with language-precision criteria that assess clarity, specificity, and terminological consistency according to recognized guidelines for accurate and unbiased scholarly expression Boutron et al. (2010); International Committee of Medical Journal Editors (2025); University of Chicago Press (2024); Garner (2019, 2013). Paragraph Quality reflects cohesion-assessment rules that examine whether individual paragraphs follow a coherent internal logic—anchoring topic statements, supporting evidence, and transitional structure—in line with authoritative reporting and specification frameworks Bastian and Moher (2021); IEEE Computer Society (2018); Harvard Law Review Association (2020); American Sociological Association (2022). Readability corresponds to accessibility-oriented requirements that evaluate whether the narrative facilitates comprehension through appropriate signaling, explanatory devices, and presentation practices as outlined in major editorial and reporting guidelines Bastian and Moher (2021); International Committee of Medical Journal Editors (2025); Garner (2013); Global Reporting Initiative (2023).
@@
Sensitive Handling corresponds to ethical-screening requirements that evaluate whether sensitive domains are addressed with neutrality, respect, and responsible language in accordance with established editorial and reporting ethics standards International Committee of Medical Journal Editors (2025); Bastian and Moher (2021); Oral History Association (2024); Organization of American Historians (2018); British Psychological Society (2021); National Commission for the Protection of Human Subjects of Biomedical and Behavioral Research (1979); American Political Science Association (2012a). Safety & Impact aligns with harm-assessment provisions that examine whether potential adverse effects, misuse risks, or disproportionate impacts are identified and mitigated under recognized guidelines for responsible scientific communication International Committee of Medical Journal Editors (2025); Guyatt et al. (2013); National Research Council (2011); World Medical Association (2013); Ecological Society of America (2021). Perspective Balance reflects fairness-evaluation criteria that assess whether multiple viewpoints and counterpositions are represented without undue bias following principles of impartiality articulated in authoritative publication and reporting frameworks International Committee of Medical Journal Editors (2025); Bastian and Moher (2021); British Philosophical Association (2024); Australasian Association of Philosophy (2023); CFA Institute (2024); American Bar Association (2023); American Association for Public Opinion Research (2021).
@@
Evidence Coverage aligns with evidence-provision requirements that evaluate whether claims are accompanied by verifiable supporting sources in accordance with established evidentiary and reporting guidelines Bastian and Moher (2021); Data Citation Synthesis Group (2014); Chang et al. (2024); International Society for Stem Cell Research (2025); CFA Institute (2020). Claim Amount corresponds to content-adequacy criteria that assess whether a report supplies all necessary factual and contextual material needed to justify its reasoning under recognized completeness and transparency standards Bastian and Moher (2021); International Organization for Standardization (2011); EQUATOR Network (2025); Linguistic Society of America (2024); The Journal of Organic Chemistry (2025); U.S. Geological Survey (2024); Global Reporting Initiative (2023). Citation Amount align with source-attribution rules that examine whether in-text citations are provided at appropriate argumentative locations following authoritative norms for evidentiary traceability International Committee of Medical Journal Editors (2025); Wilkinson et al. (2016); IEEE (2025); American Chemical Society (2021); Harvard Law Review Association (2020). Reference Amount are evaluated to determine whether the number of actually used references is appropriate and whether they collectively form a well-documented, traceable, and accessible evidence base Stanford HAI (2025); Data Citation Synthesis Group (2014); American Political Science Association (2012b).
@@
Claim Factuality reflects evidence-verification provisions that assess whether factual assertions in a report correspond to verifiable sources and documented evidence traces within established evaluative frameworks Wilkinson et al. (2016); American Physical Society (2023); National Institute of Standards and Technology (2023); Navas and others (2024); National Institute of Standards and Technology (2024); U.S. Securities and Exchange Commission (2024). Citation Support corresponds to source-justification checks that evaluate whether cited references substantively support the claims they accompany according to recognized standards for evidential accountability Wilkinson et al. (2016); American Sociological Association (2018); American Historical Association (2024); Berez-Kroeker and others (2018). Reference Support aligns with source-quality and provenance requirements that examine whether referenced materials reliably and transparently support the arguments for which they are cited, in line with authoritative data-governance and reporting criteria Wilkinson et al. (2016); Chang et al. (2024); Data Citation Synthesis Group (2014); American Psychological Association (2025); International Accounting Standards Board (2024). Reference Quality evaluates whether the set of sources consists of reproducible, trustworthy, and methodologically sound information sources, ensuring that the report’s evidentiary foundation is grounded in reliable materials American Chemical Society (2021); IEEE (2025); Stanford Encyclopedia of Philosophy (2025); World Medical Association (2013); National Commission for the Protection of Human Subjects of Biomedical and Behavioral Research (1979). Reference Diversity evaluates whether the evidence base avoids excessive concentration on a small number of sources and instead maintains a balanced level of source diversity, reducing the risk of narrow, biased, or selectively framed arguments EQUATOR Network (2025); Linguistic Society of America (2024); American Association for Public Opinion Research (2021).
@@
Domain (80) Reference
AI (8) National Institute of Standards and Technology,2023; Stanford HAI,2025; Association for Computational Linguistics,2024; NeurIPS Foundation,2025; International Conference on Machine Learning,2025; International Conference on Learning Representations,2024; Journal of Machine Learning Research,2024; Springer Nature,2025
Biology (4) Bastian and Moher,2021; Boutron et al.,2010; International Society for Stem Cell Research,2025; Percie du Sert et al.,2020
Business (4) Global Reporting Initiative,2023; Garner,2013; International Organization for Standardization,2018; IFRS Foundation,2021
Chemistry (4) American Chemical Society,2021; International Union of Pure and Applied Chemistry,2007; The Journal of Organic Chemistry,2025; National Research Council,2011
Computer Science (4) International Organization for Standardization,2011; Association for Computing Machinery,2025; IEEE Computer Society,2018; Wilkinson et al.,2016
Earth & Env. Science (4) Intergovernmental Panel on Climate Change,2019; U.S. Geological Survey,2024; Eaton and others,2024; Ecological Society of America,2021
Economics (4) Chang et al.,2024; American Economic Association,2024; The Econometric Society,2024; OECD,2011
Education (4) What Works Clearinghouse,2022; American Educational Research Association,2006; Institute of Education Sciences,2022; American Educational Research Association et al.,2014
Engineering (4) IEEE Computer Society,2025; IEEE,2025; NASA,2016; IEEE Computer Society,2018
Finance (4) International Accounting Standards Board,2024; CFA Institute,2024, 2020; U.S. Securities and Exchange Commission,2024
History (4) American Historical Association,2024; University of Chicago Press,2024; Oral History Association,2024; Organization of American Historians,2018
Law (4) Harvard Law Review Association,2020; Garner,2019; U.S. Congress,2024; American Bar Association,2023
Linguistics (4) Linguistic Society of America,2024; Comrie, B. and Haspelmath, M. and Bickel, B.,2015; Berez-Kroeker and others,2018; International Phonetic Association,1999
Mathematics (4) European Mathematical Society,2025; American Mathematical Society,2022; Society for Industrial and Applied Mathematics,2024; National Institute of Standards and Technology,2024
Medicine (4) World Medical Association,2013; von Elm and others,2007; Gagnier and others,2013; Chan and others,2013
Philosophy (4) American Philosophical Association,2024; Stanford Encyclopedia of Philosophy,2025; British Philosophical Association,2024; Australasian Association of Philosophy,2023
Physics (4) American Physical Society,2023; International Union of Pure and Applied Physics,2010; Navas and others,2024; Mohr and others,2024
Political Science (4) American Political Science Association,2012b, 2018; OECD,2020; American Political Science Association,2012a
Psychology (4) American Psychological Association,2025; Center for Open Science,2024; British Psychological Society,2021; American Educational Research Association et al.,2014
Sociology (4) American Association for Public Opinion Research,2021; Tong et al.,2007; National Commission for the Protection of Human Subjects of Biomedical and Behavioral Research,1979; American Sociological Association,2022
@@ Appendix E Rubric Structure
Dim / Sub-dim Criterion / Rubric items Description
2. Analytical Soundness └ 2.2 Reasoning 2.2.5 Criterion All major claims follow logically from the previously presented facts, data, interpretations, and assumptions, without skipped steps or unsupported leaps. [2.2.5.1 Coverage] All claims that require logical support are explicitly linked to the relevant facts, data, and interpretations (including the key evidence specified in the Expert Evaluation Guidance, EG), and no central claim is left without the necessary supporting evidence. [2.2.5.2 Quality] Each major claim is explained through clear, well-structured reasoning that shows how the underlying facts, data, and interpretations lead to that conclusion, making the logical connection explicit and easy for an expert reader to follow.
Table 11: Example Coverage/Quality rubric factors for a reasoning-focused element under 2. Analytical Soundness / 2.2 Reasoning.
@@ E.1 Hierarchical Levels
Level 2: sub-dimensions (25)
Detailed criteria that decompose each dimension into finer evaluation units; 2–5 sub-dimensions are defined under each dimension. This forms an intermediate-level unit that specifies which aspects to examine for each dimension.
@@
Thus, with 7 dimensions, 25 sub-dimensions, and 46 elements as the framework, each element is divided into Coverage/Quality factors to constitute a total of 101 specific check items. Table 11 shows an example of one element and its corresponding Coverage/Quality factors in the 2. Analytical Soundness / 2.2 Reasoning dimension.